### v2

Abstract
Context. Small bodies in Earth-like orbits, the Arjunas, are good targets for scientific exploration and mining studies as they enable low-cost missions. The subset of such objects that experience recurrent temporarily captured flyby or orbiter events, also called mini-moon episodes, are among the best ranked in terms of accessibility. Only a handful of objects are known to have engaged in such a dynamical behavior. Finding and characterizing more of them may help to expand scientific and commercial research activities in space over the next few decades. Asteroid 2024 PT5 was found recently and belongs to this group of interesting objects.
Aims. Here, we investigate the orbital context of 2024 PT5, and its spectral and rotational properties.
Methods. We studied the short-term orbital evolution of 2024 PT5 using direct -body simulations. We identified its spectral class from the visible reflectance spectrum and used photometric observations to derive its rotational properties. Observational data were obtained with the OSIRIS camera spectrograph at the 10.4 m Gran Telescopio Canarias and the Two-meter Twin Telescope.
Results. Asteroid 2024 PT5 experiences recurrent co-orbital engagements and mini-moon events of the temporarily captured flyby type. Its visible spectrum is consistent with that of an Sv-type asteroid or perhaps lunar ejecta. Its rotational period could be 1 h.
Conclusions. The discovery of 2024 PT5 confirms that temporarily captured flybys are relatively frequent and involve objects larger than a few meters, suitable as accessible targets for scientific research activities and commercial mining ventures in space.
Key Words.:
minor planets, asteroids: general – minor planets, asteroids: individual: 2024 PT5 – techniques: spectroscopic – methods: numerical – celestial mechanics1 Introduction
Both ESA and NASA are pushing into the commercial space sector supporting a nascent global space economy (see, e.g., Rausser et al. 2023; Paravano et al. 2024). Scientific exploration and mining of asteroids are activities well suited for development within the context of this fast-evolving sector (see, e.g., Sommariva 2015; Sachdeva 2018; Xie et al. 2021), but only if there are enough targets for low-cost missions. NASA’s Near-Earth Object Human Space Flight Accessible Targets Study (NHATS, Abell et al. 2012)111https://cneos.jpl.nasa.gov/nhats/ is compiling a list of near-Earth objects (NEOs) that identifies potential targets for future exploration, but such a database must include information on the composition of the targets for better mission planning. Reflectance spectroscopy can help in finding what these more accessible targets are made of (Licandro et al. 2020; de León et al. 2021).
The NHATS list includes thousands of targets, but just a few have been characterized spectroscopically (Popescu et al. 2019). Most targets with the highest number of viable trajectories (the most accessible ones) follow Earth-like paths, they are part of a secondary asteroid belt that surrounds the orbit of the Earth–Moon system. Its members define a dynamical class, the Arjunas, that follow heliocentric trajectories with orbital periods close to one sidereal year, low eccentricity, and low inclination (Rabinowitz et al. 1993). Within the 25 most accessible targets, a handful undergo recurrent temporarily captured flyby or orbiter events, also called mini-moon episodes (Granvik et al. 2012).
The recently found asteroid 2024 PT5 is part of the group of very accessible targets (de la Fuente Marcos & de la Fuente Marcos 2024). Here, we use reflectance spectroscopy, photometry, and -body simulations to confirm its nature and dynamical context. This Letter is organized as follows. In Sect. 2, we introduce the background of our research, and present the data and tools used in our analyses. In Sect. 3, we determine if 2024 PT5 is natural or artificial and its probable origin. In Sect. 4, we discuss our results and Sect. 5 summarizes our conclusions. Appendices include supporting material.
2 Context, methods, and data
In this section, we revisit dynamical concepts that are later used in our analysis. Software tools and data are also discussed here.
2.1 Dynamics background
In this work (but see Appendix A), we use the definitions in Kary & Dones (1996), Granvik et al. (2012) and Fedorets et al. (2017). Temporarily captured natural irregular satellites of Earth must have negative geocentric energy and their geocentric distance must be under three Hill radii (the Hill radius of Earth is roughly 0.01 au); in other words, the intruding NEO must approach at close range (0.03 au) and low relative velocity (1 km s-1). The only NEOs that can regularly meet the capture conditions defined by Kary & Dones (1996) are the members of the Arjuna asteroid belt (Rabinowitz et al. 1993). Some of these objects experience resonant behavior induced by the 1:1 mean-motion resonance with Earth (see, e.g., de la Fuente Marcos & de la Fuente Marcos 2013). When subjected to this resonance, the relative mean longitude of the objects with respect to our planet () oscillates about a fixed value (Morais & Morbidelli 2002) and they become Earth co-orbitals as the orbital period of the NEO closely matches that of Earth. One of the resonant states compatible with co-orbital behavior drives objects to a horseshoe trajectory with respect to Earth when the orbit is plotted in a heliocentric frame of reference rotating with our planet because in this case the value of oscillates about 180°, with an amplitude 240° (Murray & Dermott 1999). In certain circumstances, perturbed horseshoe paths may lead to temporarily captured flybys as in the case of 2022 NX1 (de la Fuente Marcos & de la Fuente Marcos 2022; de la Fuente Marcos et al. 2023) or orbiters (see Appendix B).
2.2 Data, data sources, and tools
Object A119q0V was initially reported on August 7, 2024, by the Asteroid Terrestrial-impact Last Alert System (ATLAS, Tonry et al. 2018) observing from Sutherland in South Africa. On August 14, it was announced with the provisional designation 2024 PT5 (Tonry et al. 2024). Its orbit determination in Table 1 is currently based on 247 observations with a data-arc span of 73 days, it is referred to epoch JD 2460600.5 TDB, which is the origin of time in the calculations, and corresponds to that of an Apollo-class asteroid. It was retrieved from Jet Propulsion Laboratory’s (JPL) Small-Body Database (SBDB)222https://ssd.jpl.nasa.gov/tools/sbdb_lookup.html#/ provided by the Solar System Dynamics Group (SSDG, Giorgini 2011, 2015).333https://ssd.jpl.nasa.gov/ The object attracted our attention because it approached Earth at close range and low relative velocity.
| Orbital parameter | value1 uncertainty | |
| Semimajor axis, (au) | = | 1.012305040.00000003 |
| Eccentricity, | = | 0.021476730.00000002 |
| Inclination, (°) | = | 1.5205220.000009 |
| Longitude of the ascending node, (°) | = | 305.57240.0002 |
| Argument of perihelion, (°) | = | 116.24840.0002 |
| Mean anomaly, (°) | = | 323.677200.00007 |
| Perihelion distance, (au) | = | 0.990564030.00000003 |
| Aphelion distance, (au) | = | 1.034046040.00000003 |
| Absolute magnitude, (mag) | = | 27.40.5 |
The orbit determination of 2024 PT5 resembles that of 2022 NX1, a confirmed natural object that experiences recurrent resonant co-orbital episodes and mini-moon events of the temporarily captured flyby type. Asteroid 2022 NX1 also undergoes periodic close encounters with the Earth–Moon system that make the reconstruction of its past orbital evolution and the prediction of its future behavior beyond a few decades difficult (de la Fuente Marcos & de la Fuente Marcos 2022; de la Fuente Marcos et al. 2023). In such cases, the orbital evolution has to be studied numerically considering the uncertainties of the orbit determination. The -body simulations needed to investigate the evolution of 2024 PT5 were carried out using a direct -body code described by Aarseth (2003) that is publicly available from the web site of the Institute of Astronomy of the University of Cambridge.555http://www.ast.cam.ac.uk/sverre/web/pages/nbody.htm This software implements the Hermite numerical integration scheme developed by Makino (1991). Technical details and relevant results from this code were discussed in de la Fuente Marcos & de la Fuente Marcos (2012). Our physical model included the perturbations by the eight major planets, the Moon, the barycenter of the Pluto-Charon system, and the 19 largest asteroids, Ceres, Pallas, Vesta, Hygiea, Euphrosyne, Interamnia, Davida, Herculina, Eunomia, Juno, Psyche, Europa, Thisbe, Iris, Egeria, Diotima, Amphitrite, Sylvia, and Doris. For accurate initial positions and velocities (see Appendix C), we used data from JPL’s SSDG Horizons on-line Solar system data and ephemeris computation service,666https://ssd.jpl.nasa.gov/horizons/ that are based on the DE440/441 planetary ephemeris (Park et al. 2021). Most input data were retrieved from JPL’s SBDB and Horizons using tools provided by the Python package Astroquery (Ginsburg et al. 2019) and its HorizonsClass class.777https://astroquery.readthedocs.io/en/latest/jplhorizons/jplhorizons.html
3 Results
Here, we use -body simulations to assess the current dynamical status of 2024 PT5, reflectance spectroscopy to determine its physical nature, and photometry to study its rotational state.
3.1 Orbital evolution
Figure 1 summarizes the results of our calculations, showing the time evolution of relevant parameters of 2024 PT5: the relative mean longitude, and the geocentric energy and distance. The figure shows the evolution of the nominal orbit and those of control orbits or clones with state vectors (Cartesian coordinates and velocities, see Table 2 in Appendix C) well away from the nominal one, up to from the nominal orbit determination in Table 1. The right panels in Fig. 1 show that the short-term evolution of all the control orbits matches that of the nominal orbit. Therefore, we confirm the results in de la Fuente Marcos & de la Fuente Marcos (2024) obtained using an early orbit determination. However, the left panel in Fig. 1 (see also Fig. 10) shows that recovering the past orbital evolution of 2024 PT5 prior to 1937 is difficult because a close encounter with Earth drives originally close orbits to diverge. A similar situation is found when predicting the dynamical behavior of this object beyond 2085. In other words, the orbital evolution over the time interval (87, 61) yr on the left panel of Fig. 1 can be computed precisely but beyond this time interval the current orbit determination cannot provide reliable ephemerides.
Figure 1, right panels, show that from 2023 until 2025, 2024 PT5 followed a horseshoe-like trajectory (top panel) not too different from that of 2022 NX1 in fig. 3 of de la Fuente Marcos et al. (2023). It is also observed that in 2024, between September 29 and November 25, the geocentric energy of all the control orbits had a negative value (Fig. 1, middle-right panel), and that its geocentric distance remained under 0.03 au during the same time interval (bottom panel). According to Kary & Dones (1996), 2024 PT5 becomes a temporary satellite of Earth during that period. Furthermore, as the asteroid fails to complete one revolution around Earth and following Granvik et al. (2012), 2024 PT5 undergoes a temporarily captured flyby. This result is statistically robust and consistent across the control orbits tested (104). Figure 1, left panels, show that 2024 PT5 might have experienced 1:1 resonant behavior of the horseshoe type in the past and perhaps it will show it again in the future (top panel), although the evolution is seriously affected by orbital chaos induced by close encounters with the Earth–Moon system. The middle and bottom panels show that recurrent temporarily captured flyby episodes could be possible. In fact, this object may experience further temporarily captured flybys in the future.
3.2 Spectroscopy
Three visible spectra of 2024 PT5 were obtained on the night of September 7, 2024, starting at 22:45 UTC, and using the OSIRIS camera spectrograph (Cepa et al. 2000; Cepa 2010). The instrument is installed at the 10.4 m Gran Telescopio Canarias (GTC), located at the El Roque de los Muchachos Observatory, in La Palma. It has a blue-sensitive 40964096 pixel CCD that yields a 7.8′7.8′ field of view. Observations were done using a slit width of 1.2″ (oriented to the parallactic angle) and the R300R grism (0.48–0.92 m). We obtained three individual spectra of 900 s of exposure time each, at an airmass of 1.45. Two solar analog stars (Landolt SA 112-1333 and SA115-271) were also observed with the same instrumental configuration and at a similar airmass to that of the asteroid, in order to get its reflectance spectrum. Both asteroid and solar analogs 2D spectral images were bias and flat-field corrected, background subtracted and collapsed to 1D, using an aperture equal to the distance from the center of the spatial profile to the pixel with 10% peak intensity. One-dimensional spectra were then wavelength calibrated with Xe+Ne+HgAr arc lamps, and a final, averaged spectrum for the asteroid and for each star were obtained. As a last step, we divided the spectrum of the asteroid by the spectrum of each solar analog star, and averaged the two resulting ratios. The final reflectance spectrum of 2024 PT5 is shown in Fig. 2, in green.
We used the M4AST888https://spectre.imcce.fr/m4ast/index.php/index/home online tool (Popescu et al. 2012) to get the most representative taxonomical type for 2024 PT5. This tool fits a curve to the spectrum and compares it to the ‘taxons’ defined by DeMeo et al. 2009 using a procedure. The best fit corresponds to an Sv-type (in orange in Fig. 2), a class that serves as a link between S-types (mixtures of silicates and metal) and V-types, indicative of basaltic surfaces, and with asteroid (4) Vesta as the main representative of the class. The M4AST tool was also used to compare the spectrum of the asteroid with more than 15 000 spectra of meteorites, terrestrial rocks, and lunar soils contained in the RELAB999https://sites.brown.edu/relab/relab-spectral-database/ database. Figure 2 shows the best fit in red, corresponding to a powder sample (grain sizes from 10 to 45 m) from a mare breccia of the Moon, collected by the Luna 24 mission. Interestingly, and in order of increasing values of , this best fit is followed by several samples from Moon maria collected at the Apollo 12 landing site, and a sample of a terrestrial basaltic rock (Brockenheim, Germany).
Sharkey et al. 2021 found that Earth’s quasi-satellite (469219) 2016 HO3 Kamo’oalewa (de la Fuente Marcos & de la Fuente Marcos 2016) presents a visible spectrum consistent with S- and L-type taxonomic classes, but when extended to the near-infrared, it shows a very steep spectral slope, consistent with space-weathered, lunar-like silicates. The visible spectrum of Kamo’oalewa (in light-blue), shown in Fig. 3 and digitized from Sharkey et al. 2021, is similar to that of 2022 NX1 as published in de la Fuente Marcos et al. (2023) and classified as a K-type, although we would need the near-infrared to confirm such an agreement. Note that the visible spectrum of 2024 PT5 shows a deeper absorption band at 1 m, compared to those from Kamo’oalewa and 2022 NX1, suggesting slightly less weathered silicates for the former.
3.3 Light curve
The OSIRIS instrument was also used to study the rotational properties of the object. On September 28, 2024, a series of continuous 30 s exposures were taken during one hour using the Sloan- filter and the standard CCD configuration ( binning, plate scale of 0.254 ″ pixel-1, readout time of 21 s). At the time of observation, the object had an apparent magnitude and a total apparent angular rate of 1.2 ″ min-1. Standard bias and flat-field corrections were applied to reduce the images.
Relative aperture photometry was performed on the images using three field stars as reference stars. Different aperture radii were tested, including both fixed and variable apertures with aperture corrections. The best result, obtained using a fixed aperture radius of about FWHM, is shown in Fig. 4.
The signal-to-noise ratio (SNR) of the light curve is not sufficient to unambiguously determine a rotation period. However, a variability with an amplitude of at least 0.5 mag is observed during the one-hour observing block, which is consistent with the uncertainty in in Table 1. Therefore, we consider that the object may have a rotation period of around one hour or less. This would be similar to that of Kamo’oalewa, which has an estimated rotation period of min (Sharkey et al. 2021). For an object with an absolute magnitude of 27.4, the possibility of a higher rotational frequency or even tumbling motion cannot be excluded. However, observations with higher time resolution and better SNR are required to confirm any of such behaviors.
4 Discussion
Our spectroscopic results on the surface composition of 2024 PT5 are not conclusive due to the lack of a near-infrared reflectance spectra, but they are suggestive of a Lunar origin. Within the context of the Mission Accessible Near-Earth Object Survey (MANOS, Kareta et al. 2022), Kareta et al. used the Lowell Discovery Telescope (LDT) and the NASA Infrared Telescope Facility (IRTF) to obtain the visible and near-infrared spectrum of 2024 PT5 to find that it is well matched by samples of the Moon.101010https://iawn.net/documents/20240917_19th_virtual/IAWN_2024_Kareta.pdf As pointed out above, the available data indicate that (469219) 2016 HO3 Kamo’oalewa, 2022 NX1, and 2024 PT5 may all have a common origin as Lunar ejecta. The topic of the dynamics of Lunar ejecta in the neighborhood of the Earth–Moon system was first studied by Warren (1994) and Gladman et al. (1995) arriving at the conclusion that these materials can be scattered out of their source region relatively quickly, in an effective timescale of 105 yr. This subject is now receiving renewed attention (Castro-Cisneros et al. 2023; Jiao et al. 2024) after the confirmation of Kamo’oalewa as Lunar ejecta, but it is also being linked to the origin of the NEOs involved in temporary capture events (Jedicke et al. 2024). Nevertheless, the theoretical expectations drawn from the literature indicate that Kamo’oalewa, 2022 NX1, and 2024 PT5 may have all emerged out of lunar impact craters formed during the last Myr or so.
5 Summary and conclusions
In this Letter, we presented a detailed analysis of a mini-moon engagement of the temporarily captured flyby type experienced by Arjuna asteroid 2024 PT5 in 2024, from September 29 to November 25. The study of the short-term orbital evolution of 2024 PT5 was carried out using direct -body simulations. We derived its spectral class from a visible reflectance spectrum obtained with the OSIRIS camera spectrograph at the 10.4 m Gran Telescopio Canarias. Data from the same instrument were used to perform a preliminary assessment of its rotational state. Astrometry was obtained using the Two-meter Twin Telescope (see Appendix D). Our conclusions can be summarized as follows.
1.
We confirm that 2024 PT5 is a natural object with a visible spectrum consistent with that of an Sv-type asteroid, although it could also be classified as a Lunar mare breccia.
2.
We estimate a rotational period 1 h, tumbling not excluded.
3.
We confirm that 2024 PT5 experienced a temporarily captured flyby between September 29 and November 25, 2024.
4.
The orbital evolutions of 2022 NX1 and 2024 PT5 are very similar as well as the circumstances of their temporarily captured flyby episodes.
Both ESA and NASA now emphasize a low-cost approach to NEO missions for small body science and planetary defense, including reuse and repurposing of already active missions (Freeman et al. 2024). The discovery and characterization of accessible objects like 2022 NX1 and 2024 PT5 using ground-based facilities is a necessary step before deciding if an ongoing mission can be extended to study a newly found NEO or perhaps a new low-cost mission is preferable.
Acknowledgements.
RdlFM and CdlFM thank S. J. Aarseth for providing one of the codes used in this research and A. I. Gómez de Castro for providing access to computing facilities. JdL and JL acknowledge support from the Spanish ‘Agencia Estatal de Investigación del Ministerio de Ciencia e Innovación’ (AEI-MCINN) under the grant PID2020-120464GB-I00. This work was partially supported by the Spanish ‘Agencia Estatal de Investigación (Ministerio de Ciencia e Innovación)’ under grant PID2020-116726RB-I00 /AEI/10.13039/501100011033. Based on observations made with the Gran Telescopio Canarias (GTC), installed at the Spanish Observatorio del Roque de los Muchachos of the Instituto de Astrofísica de Canarias, on the island of La Palma. This work is partly based on data obtained with the instrument OSIRIS, built by a Consortium led by the Instituto de Astrofísica de Canarias in collaboration with the Instituto de Astronomía of the Universidad Nacional Autónoma de Mexico. OSIRIS was funded by GRANTECAN and the National Plan of Astronomy and Astrophysics of the Spanish Government. This Letter includes observations made in the Two-meter Twin Telescope (TTT) and Transient Survey Telescope (TST) in the Teide Observatory of the IAC, that Light Bridges operates in the Island of Tenerife, Canary Islands (Spain). The Observing Time Rights (DTO) used for this research were provided by PEI project SIDERA24. This Letter used flash storage and GPU computing resources as Indefeasible Computer Rights (ICRs) being commissioned at the ASTRO POC project that Light Bridges will operate in the Island of Tenerife, Canary Islands (Spain). The ICRs used for this research were provided by Light Bridges in cooperation with Hewlett Packard Enterprise (HPE) and VAST DAT. This research utilizes spectra acquired by Carle M. Pieters with the NASA RELAB facility at Brown University. In preparation of this Letter, we made use of the NASA Astrophysics Data System, the ASTRO-PH e-print server, and the MPC data server.References
- Aarseth (2003) Aarseth, S. J. 2003, Gravitational N-Body Simulations (Cambridge: Cambridge University Press), 27
- Abell et al. (2012) Abell, P. A., Barbee, B. W., Mink, R. G., et al. 2012, 43rd Annual Lunar and Planetary Science Conference, 2842
- Alarcon et al. (2023) Alarcon, M. R., Licandro, J., Serra-Ricart, M., et al. 2023, PASP, 135, 055001
- Bolin et al. (2020) Bolin, B. T., Fremling, C., Holt, T. R., et al. 2020, ApJ, 900, L45
- Carusi & Valsecchi (1979) Carusi, A. & Valsecchi, G. B. 1979, Riunione della Societa Astronomica Italiana, 22, 181
- Castro-Cisneros et al. (2023) Castro-Cisneros, J. D., Malhotra, R., & Rosengren, A. J. 2023, Communications Earth and Environment, 4, 372
- Cepa et al. (2000) Cepa, J., Aguiar, M., Escalera, V. G., et al. 2000, Proc. SPIE, 4008, 623
- Cepa (2010) Cepa, J. 2010, Astrophysics and Space Science Proceedings, 14, 15
- de la Fuente Marcos & de la Fuente Marcos (2012) de la Fuente Marcos, C. & de la Fuente Marcos, R. 2012, MNRAS, 427, 728
- de la Fuente Marcos & de la Fuente Marcos (2013) de la Fuente Marcos, C. & de la Fuente Marcos, R. 2013, MNRAS, 434, L1
- de la Fuente Marcos & de la Fuente Marcos (2016) de la Fuente Marcos, C. & de la Fuente Marcos, R. 2016, MNRAS, 462, 3441
- de la Fuente Marcos & de la Fuente Marcos (2020) de la Fuente Marcos, C. & de la Fuente Marcos, R. 2020, MNRAS, 494, 1089
- de la Fuente Marcos & de la Fuente Marcos (2022) de la Fuente Marcos, C. & de la Fuente Marcos, R. 2022, RNAAS, 6, 160
- de la Fuente Marcos & de la Fuente Marcos (2024) de la Fuente Marcos, C. & de la Fuente Marcos, R. 2024, RNAAS, 8, 224
- de la Fuente Marcos et al. (2023) de la Fuente Marcos, R., de León, J., de la Fuente Marcos, C., et al. 2023, A&A, 670, L10
- de León et al. (2021) de León, J., Licandro, J., Popescu, M., et al. 2021, 7th IAA Planetary Defense Conference, 25
- DeMeo et al. (2009) DeMeo, F., Binzel, R. P., Slivan, S. M., et al. 2009, Icarus, 202, 160
- Fedorets et al. (2017) Fedorets, G., Granvik, M., & Jedicke, R. 2017, Icarus, 285, 83
- Freeman et al. (2024) Freeman, A., Fesq, L., Matousek, S., et al. 2024, Acta Astronautica, 224, 122
- Ginsburg et al. (2019) Ginsburg, A., Sipőcz, B. M., Brasseur, C. E., et al. 2019, AJ, 157, 98
- Giorgini (2011) Giorgini, J. 2011, in Journées Systèmes de Référence Spatio-temporels 2010, ed. N. Capitaine, 87–87
- Giorgini (2015) Giorgini, J. D. 2015, IAUGA, 22, 2256293
- Gladman et al. (1995) Gladman, B. J., Burns, J. A., Duncan, M. J., et al. 1995, Icarus, 118, 302
- Granvik et al. (2012) Granvik, M., Vaubaillon, J., & Jedicke, R. 2012, Icarus, 218, 262
- Jedicke et al. (2024) Jedicke, R., Alessi, E., Wiedner, N., et al. 2024, American Astronomical Society Meeting Abstracts, 243, 363.03
- Jiao et al. (2024) Jiao, Y., Cheng, B., Huang, Y., et al. 2024, Nature Astronomy, 8, 819
- Kareta et al. (2022) Kareta, T., Moskovitz, N., & Sharkey, B. 2022, AAS/DPS Meeting Abstracts, 54, 514.06
- Kary & Dones (1996) Kary, D. M. & Dones, L. 1996, Icarus, 121, 207
- Kwiatkowski et al. (2009) Kwiatkowski, T., Kryszczyńska, A., Polińska, M., et al. 2009, A&A, 495, 967
- Licandro et al. (2020) Licandro, J., Popescu, M., Oscoz, A., et al. 2020, 14th European Planetary Science Congress, 203
- Makino (1991) Makino, J. 1991, ApJ, 369, 200
- Morais & Morbidelli (2002) Morais, M. H. M. & Morbidelli, A. 2002, Icarus, 160, 1
- Murray & Dermott (1999) Murray, C. D., & Dermott, S. F. 1999, Solar System Dynamics (Cambridge: Cambridge University Press)
- Paravano et al. (2024) Paravano, A., Patrizi, M., Razzano, E., et al. 2024, Acta Astronautica, 222, 162
- Park et al. (2021) Park, R. S., Folkner, W. M., Williams, J. G., et al. 2021, AJ, 161, 105
- Popescu et al. (2012) Popescu, M., Birlan, M., & Nedelcu, D. A. 2012, A&A, 544, A130
- Popescu et al. (2019) Popescu, M., Vaduvescu, O., de León, J., et al. 2019, A&A, 627, A124
- Rabinowitz et al. (1993) Rabinowitz, D. L., Gehrels, T., Scotti, J. V., et al. 1993, Nature, 363, 704
- Rausser et al. (2023) Rausser, G., Choi, E., & Bayen, A. 2023, Proceedings of the National Academy of Science, 120, e2222013120
- Rickman & Malmort (1981) Rickman, H. & Malmort, A. M. 1981, A&A, 102, 165
- Sachdeva (2018) Sachdeva, G. S. 2018, Astropolitics, 16, 202
- Sharkey et al. (2021) Sharkey, B. N. L., Reddy, V., Malhotra, R., et al. 2021, Communications Earth and Environment, 2, 231
- Sommariva (2015) Sommariva, A. 2015, Astropolitics, 13, 25
- Tonry et al. (2018) Tonry, J. L., Denneau, L., Heinze, A. N., et al. 2018, PASP, 130, 064505
- Tonry et al. (2024) Tonry, J., Robinson, J., Fitzsimmons, A., et al. 2024, Minor Planet Electronic Circulars, 2024-P170
- Warren (1994) Warren, P. H. 1994, Icarus, 111, 338
- Xie et al. (2021) Xie, R., Bennett, N. J., & Dempster, A. G. 2021, Acta Astronautica, 181, 249
Appendix A One Hill radius versus three Hill radii
A gravitationally bound two-body system has negative total energy. Although within the two-body problem this criterion is unambiguous, subsystems within complex -body systems may require additional constraints to ensure that a certain subsystem is indeed bound. In the particular case of planets and captured objects, Carusi & Valsecchi (1979) advocated the use of a simple criterion: the planetocentric energy of the object must be negative. Rickman & Malmort (1981) recommended adding, as second restriction, that the object completes at least one revolution around the planet. Kary & Dones (1996) recognized the critical role played by the relative distance during temporary captures by planets hosted by stars and defined temporary satellite capture by a planet using two conditions that must be met at the same time: the planetocentric energy of the object must be negative and its planetocentric distance must be under three Hill radii for the planet involved. The value of the Hill radius of Earth is roughly 0.01 au so the second condition for our planet implies flybys closer than 0.03 au. This criterion involving both geocentric energy and separation was used by Granvik et al. (2012) and Fedorets et al. (2017) to define what temporarily captured natural irregular satellites of Earth are. From there, they designated captured objects that do not complete at least one revolution around Earth as temporarily captured flybys and those completing one or more revolutions as temporarily-captured orbiters. In this work, we use the definitions in Kary & Dones (1996), Granvik et al. (2012) and Fedorets et al. (2017). For a temporary capture to take place, the intruding NEO must approach at close range (0.03 au) and low relative velocity (1 km s-1).
The Hill radius is often used to define the sphere of influence of a planet within the context of the two-body problem and the restricted three-body problem (Murray & Dermott 1999). It is sometimes argued that true temporary captures can only take place when the planetocentric energy becomes negative inside the Hill radius of the planet. We believe that the restrictive use of the one Hill radius versus the three Hill radii criterion may lead to inconsistent interpretations of true temporary orbital status of these objects. In fact, if calculations are repeated giving negligible mass to the Earth–Moon system no looping around Earth is observed; in other words, the gravitational influence of Earth in these cases is not negligible beyond one Hill radius. In addition, the strict one revolution around Earth condition cited by some is often difficult to assess because the geocentric loops traveled during temporary captures are far from closed elliptical orbits.
Appendix B Mini-moons: the tally so far
So far, temporary captures by Earth, meeting the conditions in Kary & Dones (1996), Granvik et al. (2012) and Fedorets et al. (2017), have been reported for Apollo-class asteroids 2006 RH120 (Kwiatkowski et al. 2009), 2020 CD3 (Bolin et al. 2020; de la Fuente Marcos & de la Fuente Marcos 2020), 2022 NX1 (de la Fuente Marcos & de la Fuente Marcos 2022; de la Fuente Marcos et al. 2023), and 2024 PT5 (de la Fuente Marcos & de la Fuente Marcos 2024). The condition requiring negative planetocentric energy during capture is equivalent to having planetocentric eccentricity 1. Figures 5 to 8 are similar to Fig. 9 but they show the ephemerides retrieved from JPL’s Horizons instead of displaying the results of our computer simulations. They clearly show the differences between temporarily-captured orbiters (2006 RH120 and 2020 CD3) and temporarily captured flybys (2022 NX1 and 2024 PT5), but they also show that the strict application of the one Hill radius criterion may lead us to reclassify the capture episodes of 2006 RH120 and 2020 CD3 as a sequence of temporarily captured flybys, which is obviously incorrect.
Appendix C Input data and uncertainties
Here, we include the barycentric Cartesian state vector of NEO 2024 PT5. This vector and its uncertainties have been used to perform the calculations discussed above and to generate the figures that display the time evolution of the critical angle, , and the geocentric energy and distance. As an example, a new value of the -component of the state vector was computed using , where is an univariate Gaussian random number, and and are the mean value and its 1 uncertainty in Table 2.
| Component | value1 uncertainty | |
| (au) | = | 9.0234018277509992.42255220 |
| (au) | = | 4.0001870467248861.26916846 |
| (au) | = | 2.6053151788957561.43228523 |
| (au/d) | = | 7.2758556060015816.00661110 |
| (au/d) | = | 1.5789959071111641.62598869 |
| (au/d) | = | 8.6555139017843991.84634694 |
Figure 9 shows the capture episode for the nominal orbit in greater detail. The minimum value of the geocentric energy coincides with a local maximum in the geocentric distance. The capture takes place when the asteroid approaches the “apogee” of its highly asymmetric geocentric trajectory and its geocentric velocity reaches a minimum.
Figure 10 is analogous to Fig. 1 but corresponds to control orbits with state vectors separated , , and from that of the nominal orbit. Results are consistent with those presented in Sect. 3.1, the orbital evolution of 2024 PT5 over the time interval (87, 61) yr is robust but outside this interval, the evolution has to be discussed in statistical terms as the evolution of arbitrarily close orbits diverges exponentially.
Appendix D TTT and TST astrometry
Astrometric observations were obtained using three telescopes located at Teide Observatory (OT, Tenerife, Canary Islands, Spain), the TTT1 and TTT2 (Two-meter Twin Telescope), and the TST (Transient Survey Telescope) as part of the key project observations during the commissioning phase. The observational circumstances are shown in Table 3. A stacked image of 2024 PT5 obtained with TTT1 is shown in Fig. 11.
TTT121212https://ttt.iac.es/ currently has two telescopes, TTT1 and TTT2. They are two Ritchey-Chrétien optical systems, with an aperture of 0.80 m, altazimuthal mount, and two f/6.85 Nasmyth foci each. TST131313https://tst.iac.es/ is an ultra-wide-field prime corrector optical system, with an aperture of 1 m, equatorial mount, and a f/1.3 primary foci.
TTT1 was equipped with a 2Kpx2Kpx Andor iKon-L 936 camera, with a back-illuminated 13.5 m pixel-1 BEX2-DD CCD sensor, resulting in a FoV of 17.3′17.3′, and a plate scale of 0.51″ pixel-1. QHY411M141414https://www.qhyccd.com/ cameras (Alarcon et al. 2023) are installed on TTT2 and TST telescopes. The QHY411M camera sports a scientific Complementary Metal-Oxide-Semiconductor (sCMOS) image sensor with 14304 px 10748 px (151 Mpx = 302 MB) of size 3.76 m pixel-1. This setup provides an effective FoV of 3425′ (with a plate scale of 0.14″ pixel-1) for TTT2 and 2.4°1.8° (0.6″ pixel-1) for TST.
All the images were taken using the filter (Lum), that covers the 0.4 to 0.7 m wavelength range. Data reduction was done using standard procedures, correcting for bias, dark and sky flat-fielding. The astrometry was extracted using the Tycho software,151515https://www.tycho-tracker.com/.com/ applying the track and stack method. To increase object SNR, stacking was done tracking on 2024 PT5 orbit using the orbital ephemeris of the object (individual images have sidereal tracking). The ephemeris were obtained from the Minor Planet Center website.161616https://www.minorplanetcenter.net/iau/MPEph/MPEph.html
| Date obs. | Obs | exp | filter | Mag | Telescope | |
| 2024-09-10 | 20:11 | 7 | 13060s | Lum | 22.40 | Ikon-TTT1 |
| 2024-09-10 | 20:10 | 7 | 16860s | Lum | 22.17 | QHY411-TTT2 |
| 2024-09-10 | 21:22 | 11 | 12360s | Lum | 22.05 | QHY411-TST |
| 2024-09-11 | 20:09 | 6 | 13060s | Lum | 21.95 | Ikon-TTT1 |
| 2024-09-11 | 21:03 | 4 | 14160s | Lum | 22.31 | QHY411-TST |
| 2024-09-12 | 20:06 | 8 | 13060s | Lum | 22.26 | Ikon-TTT1 |
| 2024-09-13 | 20:08 | 3 | 7860s | Lum | 21.63 | Ikon-TTT1 |
| 2024-09-15 | 20:02 | 3 | 7860s | Lum | 22.51 | Ikon-TTT1 |
| 2024-09-17 | 20:01 | 2 | 4860s | Lum | 22.51 | Ikon-TTT1 |
| 2024-09-20 | 20:30 | 3 | 6060s | Lum | 22.72 | Ikon-TTT1 |
| 2024-09-21 | 19:59 | 3 | 7960s | Lum | 22.77 | Ikon-TTT1 |
| 2024-09-22 | 19:54 | 2 | 5360s | Lum | 22.18 | Ikon-TTT1 |
| 2024-09-23 | 19:52 | 2 | 7860s | Lum | 23.09 | Ikon-TTT1 |
| 2024-09-24 | 19:55 | 3 | 6560s | Lum | 22.53 | Ikon-TTT1 |
| 2024-09-25 | 19:51 | 3 | 6560s | Lum | 22.09 | Ikon-TTT1 |
| 2024-09-26 | 19:51 | 6 | 14360s | Lum | 22.17 | Ikon-TTT1 |
| 2024-09-27 | 19:47 | 3 | 7860s | Lum | 22.43 | Ikon-TTT1 |
| 2024-09-27 | 19:47 | 3 | 12660s | Lum | 22.91 | QHY411-TTT2 |
| 2024-09-27 | 20:00 | 4 | 24530s | Lum | 22.66 | QHY411-TST |
| 2024-09-28 | 19:40 | 2 | 7860s | Lum | 22.23 | Ikon-TTT1 |
| 2024-09-28 | 20:02 | 4 | 8490s | Lum | 22.66 | QHY411-TST |
| 2024-09-28 | 21:24 | 6 | 6030s | 22.66 | GTC | |
| 2024-09-29 | 19:40 | 3 | 7860s | Lum | 22.8 | Ikon-TTT1 |
| 2024-09-30 | 19:37 | 3 | 7360s | Lum | 22.1 | Ikon-TTT1 |
| 2024-10-01 | 19:50 | 3 | 8760s | Lum | 22.2 | Ikon-TTT1 |
| 2024-10-04 | 19:35 | 3 | 18260s | Lum | 22.4 | Ikon-TTT1 |
| 2024-10-05 | 19:55 | 3 | 13060s | Lum | 22.23 | Ikon-TTT1 |
| 2024-10-07 | 21:03 | 2 | 2660s | Lum | 22.10 | Ikon-TTT1 |
| 2024-10-08 | 19:40 | 3 | 7160s | Lum | 22.47 | Ikon-TTT1 |
| 2024-10-09 | 19:51 | 3 | 14360s | Lum | 22.55 | Ikon-TTT1 |
| 2024-10-10 | 19:50 | 3 | 14160s | Lum | 21.50 | Ikon-TTT1 |
| 2024-10-14 | 19:42 | 2 | 13060s | Lum | 22.55 | Ikon-TTT1 |
| 2024-10-18 | 19:45 | 2 | 10460s | Lum | 22.1 | Ikon-TTT1 |
| 2024-10-19 | 19:45 | 2 | 9160s | Lum | 22.5 | Ikon-TTT1 |